# LiRA: A Multi-Agent Framework for Reliable and Readable Literature Review Generation

### v1

@@ Appendix D Citation Quality
For the references, only the abstracts were used, as doing so would replicate the process of how SMEs generally perform the same task. Furthermore, the implementation was taken from Wang et al. (2024) with some adjustments to better handle longer inputs. Specifically, assuming a claim-reference pair is too long, the reference is split into a few sections, and then the metric checks if any split supports the claim. This assumes the rest of the references are also supported by the sources as a result.
@@ Appendix E Ablation and Modification
The other experiment performed was the addition of a researcher agent. As a method of better imitating the process of literature review writing by humans, this agent takes each reference abstract and analyzes the contents based on several pre-determined questions based on existing literature (Snyder 2019). This results in a list of findings (i.e., key points, limitations) which are then utilized by the other agents in place of the original abstract.
@@ Appendix F Retrieval Setting Statistical Tests
Table 7: The rubric criteria alongside their descriptions, adapted from (Wang et al. 2024).

### v2

@@ Appendix D Citation Quality
For the references, only the abstracts were used, as doing so would replicate the process of how SMEs generally perform the same task. Furthermore, the implementation was taken from wang_autosurvey_2024 with some adjustments to better handle longer inputs. Specifically, assuming a claim-reference pair is too long, the reference is split into a few sections, and then the metric checks if any split supports the claim. This assumes the rest of the references are also supported by the sources as a result.
@@ Appendix E Ablation and Modification
The other experiment performed was the addition of a researcher agent. As a method of better imitating the process of literature review writing by humans, this agent takes each reference abstract and analyzes the contents based on several pre-determined questions based on existing literature (snyder_literature_2019). This results in a list of findings (i.e., key points, limitations) which are then utilized by the other agents in place of the original abstract.
@@ Appendix F Retrieval Setting Statistical Tests
Table 7: The rubric criteria alongside their descriptions, adapted from (wang_autosurvey_2024).